why cos of -theta = cos theta?????... can any body explain it??

why    cos  of  -theta  =  cos theta?????...       can  any body   explain  it??

Grade:11

2 Answers

vamsi krishna
8 years ago
cos(-x)=cos(x)..........because of cos is an even function i.e., f(-x)=f(x) here negative becomes positive ….….….…..all over Trigonometric ratios:
sin(-x)=-sinx
tan(-x)=-tanx
cot(-x)=-cotx
cosec(-x)=-cosecx......................these 4 are odd functions but where as..
 
cos(-x)=cosx
sec(-x)=secx.................these 2 are even functions......
if else any doubts upload question on it.
cherry
8 years ago
(cos(-x)=cos(x)..........because of cos is an even function i.e., f(-x)=f(x) here negative becomes positive ….….….…..all over Trigonometric ratios,) :- how they became positive ???
why sin ,cot,tan,cosec are remains  negitive???